What is the slope of the line that is perpendicular to the graph of y = 2/7 x + 9?
Answer by Alan3354(69443)

What is the slope of the line that is perpendicular to the graph of y = 2/7 x + 9?
----------------
The slope of y = (2/7)x + 9 is 2/7.
All lines perpendicular, an infinite # of them, have a slope that's the negative inverse = -7/2.